# Reconstruct, Inpaint, Finetune: Dynamic Novel-view Synthesis from Monocular Videos

### v1

@@ 3.3 Test-time finetuning for target domain adaptation
At test time, to reduce domain gap arising due to different scene properties (lighting, appearance, motion) we use the source test video $\mathbf{V}_{\rm src}$ to adjust the priors of CogNVS and create self-supervised finetuning pairs, $\mathcal{D}$ as described above. We therefore adapt the model weights $\theta$ on-the-fly with $M$ gradient steps with $\eta$ step size as follows,
@@ 4.1 Experimental setup
We train CogNVS on four in-the-wild video datasets, SA-V [15], TAO [14], Youtube-VOS [16], and DAVIS [87]. We sample 3000, 3000, 4000 and 100 videos respectively from each of the datasets, giving us a total training video pool of $\approx$ 10,000 videos. For pretraining, we randomly select a new subsequence of 49-frames in every epoch and construct its training pairs. For benchmarking, we follow prior work [46, 48] and use a combination of Kubric-4D, ParallelDomain-4D [48] and Dycheck [11]. These have a held-out test set of 20, 20 and 5 videos each. Note that our evaluation on Kubric-4D, ParallelDomain-4D and Dycheck is zero-shot as the datasets are not seen during training. Since the Kubric-4D and ParallelDomain-4D are synthetic, we use their groundtruth point clouds and odometry for a fair comparison to baselines. For Dycheck, we use MegaSAM for reconstruction and align the estimated point cloud with the groundtruth to solve for scale ambiguity.
@@
For Kubric-4D, we consider GCD [48] and Gen3C [50], alongside a concurrent work, TrajectoryCrafter [51]. For ParallelDomain-4D, we consider the same baselines except Gen3C, which only evaluates on Kubric-4D, as there is no open-source implementation available yet. For Dycheck, we consider recent work like Shape-of-Motion [3], MoSca [46], CAT4D [49]. Note that we do not benchmark test-time optimization approaches on Kubric-4D and ParallelDomain-4D, because their performance degrades catastrophically on novel views that are far apart from training views. For more quantitative analysis of CAT4D, see appendix.
@@
For pixel-wise photometric evaluation, we adopt the widely used PSNR, SSIM, and LPIPS family of metrics for evaluating reconstruction quality via novel-view synthesis. We additionally benchmark the generation quality with FID and KID. This is in line with the benchmarking proposed in several diffusion-based view synthesis works [48, 50, 53, 88].
@@
During pretraining, we load the official CogVideoX-5B-I2V checkpoint and fully finetune all 42 transformer blocks. We use the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\beta_{3}=0.98$, a learning rate of $52\times 10e-$ 5, and a batch size of 8 for 12,000 steps. To fit within 48GB VRAM, we employ DeepSpeed ZeRO-2 [89] to partition model states across 8 A6000 Ada GPUs in a distributed setting. Pretraining completes in approximately 3 days.
@@
Figure 4: We show a qualitative comparison with state-of-the-art approaches for dynamic novel-view synthesis on Kubric-4D (top), ParallelDomain-4D (middle) and DyCheck (bottom). Note how reconstruction alone, either by groundtruth depth, MegaSAM [10], Shape of Motion [3], or MoSca [46] cannot synthesize a complete novel view. Optimization based approaches like Shape of Motion, and MoSca, blur the dynamic regions when fitting 4D representations. CAT4D [49], whose visuals are taken from its project page due to unavailable code, struggles to generalize. TrajectoryCrafter [51] over-hallucinates the occluded regions and does not preserve geometry. GCD [48] performs well because it was trained on Kubric-4D and ParallelDomain-4D. Our method can instead produce photorealistic and 3D-consistent novel-views for the given scenes in a zero-shot manner with test-time finetuning, even starting from point cloud renders that are incomplete and noisy (e.g., from MegaSAM for DyCheck). It is consistently able to synthesize sharp dynamic objects, which the other baselines struggle with. Please see the video in the appendix.
@@ 4.2 Comparison to state-of-the-art
Table 2: Comparison to state-of-the-art for dynamic novel-view synthesis on Dycheck. First, we note that our method can be run on top of any reconstruction approach and the better the reconstruction (e.g., replacing MegaSAM with MoSca), the better the view synthesis. Second, we see that our method can achieve state-of-the-art FID / KID scores because test-time optimization approaches [3, 46, 10] result in blurry dynamic regions and cannot hallucinate new scene content, and completely feed-forward approaches [51] cannot return precise geometry. Our method instead gets the “best of both worlds”.
@@
MegaSAM [10] 12.16 0.299 0.698 239.57 0.148
Shape-of-Motion [3] 15.30 0.476 0.494 164.29 0.073
MoSca [46] 16.22 0.472 0.586 148.18 0.063
TrajCrafter [51] 12.74 0.337 0.749 140.35 0.059
@@
We first do zero-shot benchmarking of CogNVS on two synthetic datasets that come with high-fidelity dense depth and accurate camera odometry annotations. For a fair comparison to all baselines, we use the depth and poses to backproject the given scene into a canonical coordinate frame. Given this scene, we generate self-supervised pairs for test-time finetuning. Upon inference (see Tab. 1), we find that CogNVS beats prior work on photometric evaluation with PSNR, SSIM, LPIPS, even when baselines are not evaluated zero-shot (GCD is trained on Kubric-4D and ParallelDomain-4D and Gen3C is trained on Kubric-4D). In Fig. 4, we show the plausible and realistic novel-views predicted by our method on both datasets as compared to the baselines. This is quantitatively demonstrated by better FID and KID scores. A concurrent work, TrajectoryCrafter [90] performs competitively. We also evaluate the rendered visible scene structure from groundtruth depth, for establishing a lower bound on dynamic-view synthesis.
@@
We evaluate the performance of our method on a real-world dataset of casually captured iPhone videos in Tab. 7. First, note that since CogNVS can be applied on top of reconstructions from any method, we show two variants. Better initial reconstruction (in this case, with MoSca rather than MegaSAM) allows for better dynamic view synthesis. Second, of all approaches, our method produces the most visually plausible novel views, as captured by drastically better FID and KID. Third, note how TrajectoryCrafter [51], also based on video diffusion which was the second-best method on Kubric-4D, is unable to handle the distribution shift in Dycheck (shallow field-of-view, close-up videos of moving objects) and fails to generalize. Whereas our method benefits from test-time finetuning and is able to adjust to any new data-distribution at test-time. Other test-time optimization approaches (Shape-of-Motion, MoSca) do better as long as evaluation views are close to training views, because there is only one distribution they need to fit to.
@@ 4.3 Ablation studies
We also study the usefulness of the large-scale pretraining stage with 2D videos from 4 training datasets. In this case, test-time finetuning alone with a self-supervised objective, cannot pull CogNVS out of the local minima it reaches without a good initialization. This is a common failure mode of many test-time optimization approaches that overfit to the training views but default to rendering artifacts such as blurry dynamic regions [46]. We show in Tab. 3 (Row 1 vs. 3) and Fig. 5 that pretraining is essential for data-driven robustness.
@@ Appendix A Implementation Details
For Kubric-4D and ParallelDomain-4D, we follow the official GCD evaluation protocol. On DyCheck, we consistently report evaluation metrics at an image resolution of 360 (width) × 480 (height). We render dynamic Gaussian representations from Mosca and Shape-of-Motion with a black background for fair comparison. Both methods optimize camera poses using the ground-truth novel views to improve photometric metrics; we retain this step to stay consistent with their original implementation. CAT4D, although diffusion-based, fits a 4D-GS representation (with minor extensions) after synthesizing multi-view videos. When evaluating CogNVS on MegaSAM renders, we append a static background extracted from the full input video to better capture long-term context. The effectiveness of background stacking is validated in Section B. Also note that Shape-of-Motion and MoSca optimize for evaluation camera poses during evaluation using ground-truth novel view videos. Whether CAT4D adopts this step is unknown. We do not do this camera pose optimization at test-time. Since the DyCheck evaluation sequences are more than 49-frames in length, we isolate the static scene regions and stack them in 3D across time. This accumulated background is then rendered onto each frame which helps, to a large extent, “pre-inpaint” the static background regions using fused information from multiple 49-frame length sequences.
@@
Table 5: Effect of masking strategy on Kubric-4D. We study the effect of building CogNVS as an inpainting model using other masking strategies, specifically, random and tube masking [91]. We find that random masking is the least optimal as it does not mimic the test-time scenario, tube masking does better, but our structured masking strategy is the best for the proposed structured inpainting task.
@@
Figure 7: We illustrate the different masking strategies considered, as proposed by a prior work [91]. For random masking (left), the masked out patches are different in each frame of the input video. For tube masking (center), a random set of patches is masked but this set is constant across multiple frames of the video. For our structured masking (right), we derive the mask by rendering visible scene reconstruction from the novel views.
@@ Appendix B Ablation Study
In Tab. 4, we show the effect of using different sources of reconstructions on the entire Kubric-4D evaluation set. Specifically, we compare the structure and odometry from MegaSAM [10] and the synthetic depth groundtruth from Kubric-4D [48]. We find that both quantitatively and qualitatively (c.f. Fig. 6), our pipeline benefits more from better reconstructions. This is because the quality of reconstruction directly affects the input to CogNVS, and if the input point cloud is noisy (e.g., smearing at the object borders), the prediction of the novel view also becomes inaccurate.
@@
Since CogNVS is an inpainting model, we ablate different masking strategies to train CogNVS on three sequences from Kubric-4D, instead of the proposed structured masking. In Tab. 5 and Fig. 7, we use random and tube masking from VideoMAE [91] and apply them with a 50% masking ratio on the input video sequences divided into 16 $\times$ 16 patches. We find that random masking is the least optimal as it does not resemble the structured inpainting task at test-time. Tube masking is more amenable to the test-time inpainting pattern, which reflects as better photometric and generative metrics. Of all, our structured masking obtained by rendering scene reconstructions into the novel views performs the best.
@@
In addition to the metrics reported in the main paper, we also report masked photometric errors as proposed by a prior work [11]. While this metric only evaluates the visible scene content and how any view-dependent changes were handled during novel-view synthesis, it does not encourage the generation of unseen scene regions. On this metric, CogNVS performs competitively as compared to baselines.
Table 7: We report masked perceptual quality metrics as proposed by prior work [11]. This metric only evaluates the visible regions of the scene and so does not encourage generation of unseen scene components. Note that our method performs competitively as compared to the baselines which only focus on modeling the visible scene content.
@@
MegaSAM [10] 14.60 0.517 0.609
Shape-of-Motion [3] 16.47 0.639 0.409
MoSca [46] 17.82 0.635 0.507
CAT4D [49] 17.39 0.607 0.341
TrajCrafter [51] 13.60 0.518 0.663

### v2

@@ 3.3 Test-time finetuning for target domain adaptation
At test time, to reduce domain gap arising due to different scene properties (lighting, appearance, motion) we use the source test video $\mathbf{V}_{\rm src}$ to adjust the priors of CogNVS and create self-supervised finetuning pairs, $\mathcal{D}$ as described above. We therefore adapt the model weights $\theta$ on‐the‐fly with $M$ gradient steps with $\eta$ step size as follows,
@@ 4.1 Experimental setup
We train CogNVS on four in-the-wild video datasets, SA-V Ravi et al. (2024), TAO Dave et al. (2020), Youtube-VOS Xu et al. (2018), and DAVIS Perazzi et al. (2016). We sample 3000, 3000, 4000 and 100 videos respectively from each of the datasets, giving us a total training video pool of $\approx$ 10,000 videos. For pretraining, we randomly select a new subsequence of 49-frames in every epoch and construct its training pairs. For benchmarking, we follow prior work Lei et al. (2024); Van Hoorick et al. (2024) and use a combination of Kubric-4D, ParallelDomain-4D Van Hoorick et al. (2024) and Dycheck Gao et al. (2022). These have a held-out test set of 20, 20 and 5 videos each. Note that our evaluation on Kubric-4D, ParallelDomain-4D and Dycheck is zero-shot as the datasets are not seen during training. Since the Kubric-4D and ParallelDomain-4D are synthetic, we use their groundtruth point clouds and odometry for a fair comparison to baselines. For Dycheck, we use MegaSAM for reconstruction and align the estimated point cloud with the groundtruth to solve for scale ambiguity.
@@
For Kubric-4D, we consider GCD Van Hoorick et al. (2024) and Gen3C Ren et al. (2025), alongside a concurrent work, TrajectoryCrafter YU et al. (2025). For ParallelDomain-4D, we consider the same baselines except Gen3C, which only evaluates on Kubric-4D, as there is no open-source implementation available yet. For Dycheck, we consider recent work like Shape-of-Motion Wang et al. (2024), MoSca Lei et al. (2024), CAT4D Wu et al. (2024b). Note that we do not benchmark test-time optimization approaches on Kubric-4D and ParallelDomain-4D, because their performance degrades catastrophically on novel views that are far apart from training views. For more quantitative analysis of CAT4D, see appendix.
@@
For pixel-wise photometric evaluation, we adopt the widely used PSNR, SSIM, and LPIPS family of metrics for evaluating reconstruction quality via novel-view synthesis. We additionally benchmark the generation quality with FID and KID. This is in line with the benchmarking proposed in several diffusion-based view synthesis works Van Hoorick et al. (2024); Ren et al. (2025); Liu et al. (2023b); Tung et al. (2024).
@@
During pretraining, we load the official CogVideoX-5B-I2V checkpoint and fully finetune all 42 transformer blocks. We use the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\beta_{3}=0.98$, a learning rate of $2\times 10e-5$, and a batch size of 8 for 12,000 steps. To fit within 48GB VRAM, we employ DeepSpeed ZeRO-2 Rajbhandari et al. (2020) to partition model states across 8 A6000 Ada GPUs in a distributed setting. Pretraining completes in approximately 3 days.
@@
Figure 4: We show a qualitative comparison with state-of-the-art approaches for dynamic novel-view synthesis on Kubric-4D (top), ParallelDomain-4D (middle) and DyCheck (bottom). Note how reconstruction alone, either by groundtruth depth, MegaSAM Li et al. (2024b), Shape of Motion Wang et al. (2024), or MoSca Lei et al. (2024) cannot synthesize a complete novel view. Optimization based approaches like Shape of Motion, and MoSca, blur the dynamic regions when fitting 4D representations. CAT4D Wu et al. (2024b), whose visuals are taken from its project page due to unavailable code, struggles to generalize. TrajectoryCrafter YU et al. (2025) over-hallucinates the occluded regions and does not preserve geometry. GCD Van Hoorick et al. (2024) performs well because it was trained on Kubric-4D and ParallelDomain-4D. Our method can instead produce photorealistic and 3D-consistent novel-views for the given scenes in a zero-shot manner with test-time finetuning, even starting from point cloud renders that are incomplete and noisy (e.g., from MegaSAM for DyCheck). It is consistently able to synthesize sharp dynamic objects, which the other baselines struggle with. Please see the video in the appendix.
@@ 4.2 Comparison to state-of-the-art
Table 2: Comparison to state-of-the-art for dynamic novel‑view synthesis on Dycheck. First, we note that our method can be run on top of any reconstruction approach and the better the reconstruction (e.g., replacing MegaSAM with MoSca), the better the view synthesis. Second, we see that our method can achieve state-of-the-art FID / KID scores because test-time optimization approaches Wang et al. (2024); Lei et al. (2024); Li et al. (2024b) result in blurry dynamic regions and cannot hallucinate new scene content, and completely feed-forward approaches YU et al. (2025) cannot return precise geometry. Our method instead gets the “best of both worlds”.
@@
MegaSAM Li et al. (2024b) 12.16 0.299 0.698 239.57 0.148
Shape-of-Motion Wang et al. (2024) 15.30 0.476 0.494 164.29 0.073
MoSca Lei et al. (2024) 16.22 0.472 0.586 148.18 0.063
TrajCrafter YU et al. (2025) 12.74 0.337 0.749 140.35 0.059
@@
We first do zero-shot benchmarking of CogNVS on two synthetic datasets that come with high-fidelity dense depth and accurate camera odometry annotations. For a fair comparison to all baselines, we use the depth and poses to backproject the given scene into a canonical coordinate frame. Given this scene, we generate self-supervised pairs for test-time finetuning. Upon inference (see Tab. 1), we find that CogNVS beats prior work on photometric evaluation with PSNR, SSIM, LPIPS, even when baselines are not evaluated zero-shot (GCD is trained on Kubric-4D and ParallelDomain-4D and Gen3C is trained on Kubric-4D). In Fig. 4, we show the plausible and realistic novel-views predicted by our method on both datasets as compared to the baselines. This is quantitatively demonstrated by better FID and KID scores. A concurrent work, TrajectoryCrafter Xiao et al. (2025) performs competitively. We also evaluate the rendered visible scene structure from groundtruth depth, for establishing a lower bound on dynamic-view synthesis.
@@
We evaluate the performance of our method on a real-world dataset of casually captured iPhone videos in Tab. 8. First, note that since CogNVS can be applied on top of reconstructions from any method, we show two variants. Better initial reconstruction (in this case, with MoSca rather than MegaSAM) allows for better dynamic view synthesis. Second, of all approaches, our method produces the most visually plausible novel views, as captured by drastically better FID and KID. Third, note how TrajectoryCrafter YU et al. (2025), also based on video diffusion which was the second-best method on Kubric-4D, is unable to handle the distribution shift in Dycheck (shallow field-of-view, close-up videos of moving objects) and fails to generalize. Whereas our method benefits from test-time finetuning and is able to adjust to any new data-distribution at test-time. Other test-time optimization approaches (Shape-of-Motion, MoSca) do better as long as evaluation views are close to training views, because there is only one distribution they need to fit to.
Runtime analysis
We conduct a runtime analysis of CogNVS against other state-of-the-art baselines, as shown in Tab. 3, covering both stages of test-time optimization and rendering/inference. Specifically, we measure runtime by evaluating each method on the DyCheck evaluation set, with an average video length of 400 frames. Under our default hardware configuration, CogNVS requires 140 min for 400 steps of test-time fine-tuning on these long videos, followed by an additional 5 minutes for rendering. Note that this duration can be flexibly traded off with computational cost, which has recently been referred to as “inference-time scaling” in diffusion models Ma et al. (2025). For example, as shown in Fig. 8 (left) in the appendix, our method achieves 96% of its final performance with only half the fine-tuning steps.
Table 3: Runtime analysis of test-time optimization and feed-forward approaches on DyCheck. We report the time taken for optimization / test-time finetuning for all approaches in addition to the final inference / rendering duration. In general, our optimization is faster than some test-time optimization approaches with the additional benefit of being able to inpaint/hallucinate unknown regions in a spatiotemporally consistent manner. Additionally, our inference is on par with other feed-forward methods.
Method Optimization Rendering / Inference
MegaSAM Li et al. (2024b) 9 min Real-time
MoSca Lei et al. (2024) 66 min Real-time
Shape of Motion Wang et al. (2024) 237 min Real-time
GCD Van Hoorick et al. (2024) - 2 min
TrajCrafter YU et al. (2025) - 5 min
CogNVS 140 min 5 min
@@ 4.3 Ablation studies
We also study the usefulness of the large-scale pretraining stage with 2D videos from 4 training datasets. In this case, test-time finetuning alone with a self-supervised objective, cannot pull CogNVS out of the local minima it reaches without a good initialization. This is a common failure mode of many test-time optimization approaches that overfit to the training views but default to rendering artifacts such as blurry dynamic regions Lei et al. (2024). We show in Tab. 4 (Row 1 vs. 3) and Fig. 5 that pretraining is essential for data-driven robustness.
@@ 5 Discussion
Acknowledgements: This work was supported in part by Bosch and by the Intelligence Advanced Research Projects Activity (IARPA) via Department of Interior/Interior Business Center (DOI/IBC) contract number 140D0423C0074. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright annotation thereon. Disclaimer: The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of IARPA, DOI/IBC, or the U.S. Government.
@@ Appendix A Implementation Details
For Kubric-4D and ParallelDomain-4D, we follow the official GCD evaluation protocol, using the 20-sequence sub–test set and an input resolution of 576 (width) × 384 (height). On DyCheck, we follow the official Shape-of-Motion and Mosca evaluation protocols, and we consistently report evaluation metrics at an image resolution of 720 (width) × 960 (height) on the five standard test sequences: apple, block, paper-windmill, spin, and teddy. We render dynamic Gaussian representations from Mosca and Shape-of-Motion with a black background for fair comparison. Both methods optimize camera poses using the ground-truth novel views to improve photometric metrics; we retain this step to stay consistent with their original implementation. CAT4D, although diffusion-based, fits a 4D-GS representation (with minor extensions) after synthesizing multi-view videos. When evaluating CogNVS on MegaSAM renders, we append a static background extracted from the full input video to better capture long-term context. The effectiveness of background stacking is validated in Section B. Also note that Shape-of-Motion and MoSca optimize for evaluation camera poses during evaluation using ground-truth novel view videos. Whether CAT4D adopts this step is unknown. We do not do this camera pose optimization at test-time. Since the DyCheck evaluation sequences are more than 49-frames in length, we isolate the static scene regions and stack them in 3D across time. This accumulated background is then rendered onto each frame which helps, to a large extent, “pre-inpaint” the static background regions using fused information from multiple 49-frame length sequences.
@@
Table 6: Effect of masking strategy on Kubric-4D. We study the effect of building CogNVS as an inpainting model using other masking strategies, specifically, random and tube masking Tong et al. (2022). We find that random masking is the least optimal as it does not mimic the test-time scenario, tube masking does better, but our structured masking strategy is the best for the proposed structured inpainting task.
@@
Figure 7: We illustrate the different masking strategies considered, as proposed by a prior work Tong et al. (2022). For random masking (left), the masked out patches are different in each frame of the input video. For tube masking (center), a random set of patches is masked but this set is constant across multiple frames of the video. For our structured masking (right), we derive the mask by rendering visible scene reconstruction from the novel views.
@@ Appendix B Ablation Study
In Tab. 5, we show the effect of using different sources of reconstructions on the entire Kubric-4D evaluation set. Specifically, we compare the structure and odometry from MegaSAM Li et al. (2024b) and the synthetic depth groundtruth from Kubric-4D Van Hoorick et al. (2024). We find that both quantitatively and qualitatively (c.f. Fig. 6), our pipeline benefits more from better reconstructions. This is because the quality of reconstruction directly affects the input to CogNVS, and if the input point cloud is noisy (e.g., smearing at the object borders), the prediction of the novel view also becomes inaccurate.
@@
Since CogNVS is an inpainting model, we ablate different masking strategies to train CogNVS on three sequences from Kubric-4D, instead of the proposed structured masking. In Tab. 6 and Fig. 7, we use random and tube masking from VideoMAE Tong et al. (2022) and apply them with a 50% masking ratio on the input video sequences divided into 16 $\times$ 16 patches. We find that random masking is the least optimal as it does not resemble the structured inpainting task at test-time. Tube masking is more amenable to the test-time inpainting pattern, which reflects as better photometric and generative metrics. Of all, our structured masking obtained by rendering scene reconstructions into the novel views performs the best.
@@
In addition to the metrics reported in the main paper, we also report masked photometric errors as proposed by a prior work Gao et al. (2022). While this metric only evaluates the visible scene content and how any view-dependent changes were handled during novel-view synthesis, it does not encourage the generation of unseen scene regions. On this metric, CogNVS performs competitively as compared to baselines.
Maximum novel view synthesis angle
As demonstrated in Kubric 4D evaluation and several in-the-wild examples, CogNVS can plausibly generate novel views with up to 90 degrees of variation (up, down, left, or right) in general with a single forward pass. Moreover, our model can produce even more extreme viewpoints by progressively generating new views conditioned on past generations – a fairly straightforward and common strategy adopted by many generative models Höllein et al. (2023); Yu et al. (2024a); Watson et al. (2025). We illustrate such a progressively generated panoramic view in Fig. 10. Additionally, we want to point out that our method is not able to generate table top views of object-centric scenes (inward facing 360 degree view) beyond 90 degrees of camera deviation, likely because such training data was not seen by the model during pretraining.
Figure 10: We create a panoramic view of the outdoor scene through progressive generation. Specifically, we divide the 360° camera trajectory into four 90° segments of 49 frames each, reconstructing the previously generated views before reprojecting and inpainting the next 90° novel view. This qualitative analysis suggests that CogNVS can still generate 3D plausible and temporally coherent novel-views even in such extreme cases.
Table 8: We report masked perceptual quality metrics as proposed by prior work Gao et al. (2022). This metric only evaluates the visible regions of the scene and so does not encourage generation of unseen scene components. Note that our method performs competitively as compared to the baselines which only focus on modeling the visible scene content.
@@
MegaSAM Li et al. (2024b) 14.60 0.517 0.609
Shape-of-Motion Wang et al. (2024) 16.47 0.639 0.409
MoSca Lei et al. (2024) 17.82 0.635 0.507
CAT4D Wu et al. (2024b) 17.39 0.607 0.341
TrajCrafter YU et al. (2025) 13.60 0.518 0.663